the horizontal surface on which the block slides is frictionless. If F=30 N and M = 8kg, what is the magnitude of the resulting acceleration of the block
Answer #1

F=force= 30N

M= mass =8Kg

a=?

By Newton's second law

F=Ma

a = F/M = 30/8

=3.75m/sec^2